# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems: A Computational Examination of the Korteweg de-Vries and Lorenz equations

### v1

@@ Abstract.
In this work, we study the applicability of the Azouani-Olson-Titi (AOT) nudging algorithm for continuous data assimilation to evolutionary dynamical systems that are not dissipative. Specifically, we apply the AOT algorithm to the Korteweg de-Vries (KdV) equation and a partially dissipative variant of the Lorenz 1963 system. Our analysis reveals that the KdV equation lacks the finitely many determining modes property, leading to the construction of infinitely many solutions with exactly the same sparse observational data, which data assimilation methods cannot distinguish between. We numerically verify that the AOT algorithm successfully recovers these counterexamples for the damped and driven KdV equation, as studied in [55], which is dissipative. Additionally, we demonstrate numerically that the AOT algorithm is not effective in accurately recovering solutions for a partially dissipative variant of the Lorenz 1963 system.
@@ 1. Introduction
This work focuses on characterizing the class of systems for which one should not expect the implementation of the Azouani-Olson-Titi (AOT) continuous data assimilation algorithm, introduced in [4, 5]. Since its development, this algorithm has been the subject of much study, both analytically [1, 7, 8, 9, 10, 11, 16, 18, 21, 20, 27, 29, 31, 32, 33, 34, 35, 38, 39, 42, 41, 43, 51, 53, 54, 55, 60, 66, 67, 69, 71, 73, 80] and computationally [2, 19, 24, 26, 40, 44, 61, 62, 65, 59]. We note that the AOT algorithm has been referenced a few times in the literature under the names interpolated nudging or simply nudging, among others. In the context of classical methods of data assimilation, nudging refers to a method developed by Anthes and Hokes [3, 50] in the 1970’s for numerical weather prediction. The AOT algorithm is superficially similar to nudging, but is technically and conceptually different. In the classical algorithm of Anthes and Hokes one nudges the system only at the discrete nodal points of the observational data. On the other hand, the AOT algorithm works by constructing a spatial approximation interpolation operator based on the observational data and then nudges at every point in the vicinity of the observation nodal point using the interpolation operator. Nevertheless, for simplicity, we will refer to the AOT algorithm as “nudging” for the remainder of this paper. Notably, the AOT data assimilation algorithm was designed for dissipative evolutionary nonlinear systems motivated by, and capitalizing on, the fact that the long-term dynamics of such systems are determined by the dynamics of large spatial scales (see, e.g., [25] and the references therein). Such systems include for example the 2D Navier-Stokes equations, the 2D Rayleigh-Bernard convection problem, as well as the 1D Kuramoto-Sivashinsky equation, and some nonlinear reaction-diffusion systems, to name some (cf. [75]).
@@
In this work, we are concerned with the application of the nudging algorithm to systems that are not known to be dissipative. For instance, the shallow water equations have long been a simplified model for the ocean to which data assimilation algorithms have been applied. Contrary to the case of dissipative dynamical systems [25], it is unknown whether these equations have finitely many determining parameters, such as finitely many Fourier modes or nodal values which determine the solution uniquely. Yet researchers still apply data assimilation to these systems, see, e.g., [14, 58, 76]. The property, for example, of finitely many determining Fourier modes [37, 56, 25] is an essential first litmus test for data assimilation to work, as it ensures the existence of a minimum length scale at which measurements and observations determine the asymptotic behavior of the system we wish to predict. Without this property, there is no guarantee that any amount of observations will be sufficient to recover the dynamics of the true solution one wishes to predict.
Another classical example of a system that is not dissipative is the 2D damped-forced incompressible Euler equations, which have long been thought to lack the property of having finitely many determining modes. Unfortunately, we cannot even test data assimilation methods computationally on the 2D Euler equations, even though they are globally well-posed. It was shown in [30], that if we consider a given solution, $u$, to the 2D incompressible Euler equations (equipped with periodic boundary conditions) that is supported over finitely many Fourier modes for all times $t\in(0,\infty)$, then $u$ is necessarily constant in time. This indicates that solutions that are evolving in time develop arbitrarily small length scales as time advances, and thus one cannot accurately simulate the long time dynamics of the equations with a fixed spatial resolution. Moreover, utilizing adaptive mesh schemes is not enough to resolve this issue, as the computer running the simulations will eventually run out of memory and be unable to adequately refine the mesh. We encounter similar issues when even attempting to simulate the damped and driven 2D incompressible Euler system, even though it is known to have a global weak attractor, which is not necessarily finite-dimensional [47].These equations are of interest as the damping term dissipates energy and enstrophy, yet it does not introduce any smoothing effects unlike the addition of viscosity. For an in-depth look at the analytical properties of this system, see e.g. [6, 22, 23, 47, 52] and the references within.
@@
As the damped and driven Euler equations appear to be ill-suited for computational study with regard to the efficacy of data assimilation, for the remainder of this paper we will instead consider the 1D Korteweg de-Vries (KdV) equation as well as a partially dissipative variant of the Lorenz 1963 system. We consider the KdV equation in particular, as it was proven in [45, 46] that the system has a finite-dimensional global attractor in the presence of a damping and forcing terms. Moreover, it was shown in [55] that the nudging algorithm recovers the true solution given sufficiently many observed Fourier modes. Without the presence of forcing and damping, the KdV equation is a dispersive equation that has infinitely many conserved quantities, the $L^{2}$ norm being among them. The lack of dissipation and these conservation properties make this equation ideal to consider as a relatively simple 1D model to explore why data assimilation fails to recover solutions to systems without finitely many determining modes. The 1963 Lorenz system is a chaotic system that has been used to test methods of data assimilation. We note that the AOT algorithm has applied to the fully dissipative 1963 Lorenz system in [12, 72, 17].
@@
The remainder of this manuscript is organized as follows: in Section 2 we discuss various mathematical preliminaries regarding the formulation of the KdV equation and we show that, unlike the damped and driven case, the standard version of KdV does not have the finitely many determining modes property. In Section 3 we establish the numerical framework we use to investigate the efficacy of data assimilation computationally. Section 4 is split into two subsections. In Section 4.1 we show the numerical construction of pathological examples where nudging fails to recover the true solution for the KdV equations. In Section 4.2 we verify the effectiveness of nudging on the counterexamples from Section 4.1 when the system is dissipative. In Section 5 we consider a partially dissipative variant of the Lorenz system and we show that nudging fails to recover the reference solution in this setting. We end this work in Section 6 with a discussion of the conclusions we reach from our numerical results.
2. Preliminaries and Analytical Results
@@
The nudging algorithm is given as follows. Consider the following evolution system:
$u_{t}=F(u),$ (2.1)
@@
where here $F$ is a given differential operator that is potentially nonlocal and nonlinear. We will specifically be considering solutions of Equation 2.1 that lie in, e.g., $L^{2}$ subject to periodic boundary conditions.
@@
In the remainder of this section we will utilize $H^{m}_{\text{per}}$, with $m$ being a non-negative integer, to denote the closure of the set of $L$-periodic trigonometric polynomials on $\mathbf{R}$, with respect to the standard Hilbert norm:
[…]
@@
Definition 2.0.1 (Determining modes, as given in [37, 56, 25, 36]).
We say that Equation 2.1 has finitely many determining modes if there exists a positive integer $M>0$ such that for any two given solutions $u_{1}$ and $u_{2}$ of Equation 2.1 the following holds:
@@
$\lim_{t\to\infty}\norm{P_{M}(u_{1}(t)-u_{2}(t))}_{L^{2}}=0,$ (2.3)
@@
$\lim_{t\to\infty}\norm{u_{1}(t)-u_{2}(t)}_{L^{2}}=0.$ (2.4)
@@
We assume that Equation 2.1 is globally well-posed and for such systems we define the related nudged system, introduced in [4, 5], as follows:
$v_{t}=F(v)+\mu I_{h}(u-v).$ (2.5)
Here $I_{h}$ is a linear spatial interpolant with associated length scale $h$ and $\mu$ is a constant feedback-control nudging parameter. For our purposes we will consider only $I_{h}:=P_{M}$, where $P_{M}$ is given to be a projection operator onto the lowest $M$ Fourier modes, i.e. $h\sim\frac{L}{M}$, where $L$ is a typical large spatial scale for the underlying evolutionary equation Equation 2.1. As we require $P_{M}u$ to be well defined we will only consider solutions to Equation 2.1 that are in $L^{2}_{\text{per}}$ for all positive time.
@@
$u_{t}+uu_{x}+\delta^{2}u_{xxx}=0.$ (2.6)
Here $u:\mathbf{R}\times[0,\infty)\to\mathbf{R}$ represents the amplitude of a wave and $\delta$ is the dispersive coefficient. We consider the KdV equation equipped with the periodic boundary condition $u(x,t)=u(x+2,t)$, for all $x\in\mathbf{R}$ and all time $t\geq 0$ with spatial mean zero. Hence, the typical large length scale in this case is $L=2$. We note that choosing the length of the spatial domain to be 2 is arbitrary and was chosen as this choice of domain appears in the literature, see, e.g., [79].
@@
$v_{t}+vv_{x}+\delta^{2}v_{xxx}=\mu P_{M}(u-v).$ (2.7)
Here $\mu>0$ is a feedback-control nudging parameter and $P_{M}$ is given to be the projection operator onto the lowest $M$ Fourier modes. Here $u$ is the unknown reference solution of Equation 2.6, for which the first $M$ Fourier modes are observed (given).
@@
[…]
For simplicity of presentation, we omit the time index and, unless otherwise stated, we write $\hat{f}_{k}:=\hat{f}(k,t)$.
@@
Proposition 2.0.2.
Let $k\in\mathbf{Z}\setminus\{0\}$ and $\phi\in L^{2}_{\text{per}}$ with $\phi$ having spatial mean zero. The following hold true:
@@
    $\phi(x+\frac{L}{k})=\phi(x)$ for a.e. $x\in\mathbf{R}$ if and only if
    $\phi(x)=\sum_{n\in k\mathbf{Z}}\hat{\phi}_{n}e^{i\frac{2\pi}{L}nx}.$ (2.9)
@@
    Let $M$ be a positive integer and let $k$ be an integer satisfying |$\absolutevalue{k}\geq M$. Suppose $\phi(x+\frac{L}{k})=\phi(x)$ for a.e. $x\in\mathbf{R}$, then
    $P_{M}\phi=0$ (2.10)
@@
Proposition 2.0.3.
Suppose $k\in\mathbf{Z}\setminus\{0\}$ and $u^{in},f\in H^{2}_{\text{per}}$ such that $f(x+\frac{L}{k})=f(x)$ and $u^{in}(x+\frac{L}{k})=u^{in}(x)$ for all $x\in\mathbf{R}$ and $u^{in}$ and $f$ both have spatial mean zero. Let $u(x,t)$ be the corresponding solution of the damped and driven KdV equation:
$u_{t}+uu_{x}+\delta^{2}u_{xxx}+\gamma u=\beta f,$ (2.11)
$u(x,0)=u^{in}(x),$ (2.12)
with $\gamma\geq 0$ and |$\absolutevalue{\beta}\geq 0$, subject to periodic boundary conditions with period $L$.
Then $u(x+\frac{L}{k},t)=u(x,t)$ for a.e. $x\in\mathbf{R}$ and all $t\in\mathbf{R}$. Moreover, for every $0<T<\infty$, the solution $u$, restricted to the time interval $[-T,T]$, is continuous and bounded, as in [13].

### v2

@@ Abstract.
In this work, we study the applicability of the Azouani-Olson-Titi (AOT) nudging algorithm for continuous data assimilation to evolutionary dynamical systems that are not dissipative. Specifically, we apply the AOT algorithm a partially dissipative variant of the Lorenz 1963 system, the Korteweg-de Vries (KdV) in 1D, and the 2D incompressible Euler equations. Our analysis reveals that both the Euler and KdV equations lack the finitely many determining modes property, leading to the construction of infinitely many solutions with exactly the same sparse observational data, which data assimilation methods cannot distinguish between. Simultaneously, we numerically verify that the AOT algorithm successfully recovers these counterexamples for the damped and driven KdV equation, which is dissipative. Additionally, to further support our argument, we present numerical evidence showing that the AOT algorithm is ineffective at accurately recovering solutions for a partially dissipative variant of the Lorenz 1963 system.
@@ 1. Introduction
In this work, we are concerned with the application of data assimilation algorithms to systems that are not known to be dissipative. By dissipative, we mean that equations possess a compact absorbing ball or set in a relevant vector space (e.g. $L^{2}$ or $H^{1}$), which is a set that all solutions to the system eventually enter into and remain within for all future times. The existence of such an absorbing ball (or set) can be used to show the existence of a global attractor with finite fractal dimension, which in turns yields a finite number of parameters that determine the asymptotic in time behavior of solutions for dissipative dynamical systems. One relevant example of data assimilation being applied to non-dissipative dynamical systems is that of the shallow water equations, which have long been used as a simplified model for the ocean dynamics. Contrary to the case of dissipative dynamical systems [23], it is unknown whether these equations have finitely many determining parameters, such as finitely many Fourier modes or nodal values which determine the solution uniquely. Yet researchers still apply data assimilation to these systems, see, e.g., [13, 56, 77]. The property, for example, of finitely many determining Fourier modes [37, 54, 23] is an essential first litmus test for data assimilation to work, as it ensures the existence of a minimum length scale at which measurements and observations determine the asymptotic behavior of the system we wish to predict. Without this property, there is no guarantee that any amount of observations will be sufficient to recover the dynamics of the true solution one wishes to predict.
This work focuses on characterizing the class of systems for which one should not expect the implementation of the Azouani-Olson-Titi (AOT) continuous data assimilation algorithm, introduced in [4, 5], is able to effectively recover. Since its development, this algorithm has been the subject of much study, both analytically [1, 6, 7, 8, 9, 10, 16, 18, 21, 20, 26, 28, 31, 32, 33, 34, 35, 38, 39, 42, 41, 43, 50, 51, 52, 53, 58, 65, 66, 68, 70, 73, 81] and computationally [2, 19, 22, 25, 40, 44, 59, 60, 63, 57]. We note that the AOT algorithm has been referenced a few times in the literature under the names interpolated nudging or simply nudging, among others. In the context of classical methods of data assimilation, nudging refers to a method developed by Anthes and Hokes [3, 49] in the 1970’s for numerical weather prediction. The AOT algorithm is superficially similar to nudging, but is technically and conceptually different. In the classical algorithm of Anthes and Hokes one nudges the system only at the discrete nodal points of the observational data. On the other hand, the AOT algorithm works by constructing a spatial approximation interpolation operator based on the observational data and then nudges at every point in the vicinity of the observation nodal point using the interpolation operator. Nevertheless, for simplicity, we will refer to the AOT algorithm as “nudging” for the remainder of this paper. Notably, the AOT data assimilation algorithm was designed for dissipative evolutionary nonlinear systems motivated by, and capitalizing on, the fact that the long-term dynamics of such systems are determined by the dynamics of large spatial scales (see, e.g., [23] and the references therein). Such systems include for example the 2D Navier-Stokes equations, the 2D Rayleigh-Bernard convection problem, as well as the 1D Kuramoto-Sivashinsky equation, and some nonlinear reaction-diffusion systems, to name some (cf. [76]).
We note that the notion of dissipation for dynamical systems is often conflated with the inclusion of diffusion and viscous effects. We note that inclusion of a Laplacian term is one way to obtain a dissipative system as we have described, however dissipation can also be realized through damping terms, as in the Korteweg-de Vries equation (see e.g. [46, 45]). We note that for many dynamical systems relevant to geophysics, the model equations are assumed to be not be dissipative. In practice, modelers often re-introduce small artificial diffusion terms purely for numerical stability. We note however that the inclusion of such terms may guarantee that the model equations may be dissipative even though the underlying original system is not.
In this work we consider a few illustrative examples of non-dissipative dynamical systems in order to demonstrate that AOT data assimilation algorithm cannot be utilized to accurately recover the true solution. We want to illustrate that in the case of the PDEs we consider here, that the failure of the AOT algorithm, but rather is an intrinsic property of the equations considered here due to the convolutions structure of the nonlinearity. By this we mean that the pathologic solutions we consider here present an issue for any method of data assimilation, as the solutions are not uniquely identifiable from a fixed number of observed Fourier modes.
@@
In this work, we will consider three examples of non-dissipative dynamical systems; a partially dissipative variant of the 1963 Lorenz system, the classical Korteweg-de Vries equation in 1D, and the incompressible Euler equations in 2D. The 1963 Lorenz system is a chaotic system of coupled ODEs that has been used to test methods of data assimilation. We note that the AOT algorithm has applied to the fully dissipative 1963 Lorenz system in [11, 71, 17]. We consider the KdV equation in particular, as it was proven in [45, 46] that the system has a finite-dimensional global attractor in the presence of a damping and forcing terms. Moreover, it was shown in [53] that the nudging algorithm recovers the true solution given sufficiently many observed Fourier modes. Without the presence of forcing and damping, the KdV equation is a dispersive equation that has infinitely many conserved quantities, the $L^{2}$ norm being among them. The lack of dissipation and these conservation properties make this equation ideal to consider as a relatively simple 1D model to explore why data assimilation fails to recover solutions to systems without finitely many determining modes. We note that while we do consider the KdV equation, the analytical findings present here extend to a broader class of PDEs that conserve the $L^{2}$ norm featuring the same nonlinearity.
@@
The remainder of this manuscript is organized as follows: in Section˜2 we consider a partially dissipative variant of the Lorenz system and we show that nudging fails to recover the reference solution in this setting. We begin our discussion of the application of nudging to recover the solutions to the KdV equation in Section˜3.1, where we discuss various mathematical preliminaries regarding the formulation of the KdV equation and we show that, unlike the damped and driven case, the standard version of KdV does not have the finitely many determining modes property. In Section˜3.2 we show the numerical construction of pathological examples where nudging fails to recover the true solution for the KdV equations and we additionally verify the validity of the AOT nudging algorithm for these counterexamples when the system is dissipative. In Section˜4 we extend our results from the KdV regime to the case of 2D incompressible Euler equations. In Section˜4.1 we outline the analytical construction of pathological counterexamples for this regime and in Section˜4.2 we demonstrate these findings computationally. We end this work in Section˜5 with a discussion of the conclusions we reach from our numerical results.
2. Lorenz 1963 System
$\begin{cases}\dot{X}=-\sigma X+\sigma Y,&X(0)=X_{0}\\ \dot{Y}=-\sigma X-Y-XZ,&Y(0)=Y_{0}\\ \dot{Z}=-bZ+XY-b(r+\sigma),&Z(0)=Z_{0}.\end{cases}$ (1)
It is also known that applying the AOT algorithm, or a synchronization scheme, e.g., [48, 47], by observing only the $X$ component is sufficient to synchronize the nudged solution with the reference solution (see e.g. [48, 11, 47]). In particular, we can consider the following three different synchronization (nudging) schemes from [47] (which was motivated by [67]) to understand when data assimilation works for this system.
$\begin{cases}\dot{X}=-\sigma X+\sigma Y,&X(0)=X_{0}\\ \dot{y}=-\sigma X-y-Xz,&y(0)=y_{0}\\ \dot{z}=-bz+Xy-b(r+\sigma),&z(0)=z_{0}.\end{cases}$ (2)
$\begin{cases}\dot{x}=-\sigma x+\sigma Y,&x(0)=x_{0}\\ \dot{Y}=-\sigma X-Y-XZ,&Y(0)=Y_{0}\\ \dot{z}=-bz+xY-b(r+\sigma),&z(0)=z_{0}.\end{cases}$ (3)
$\begin{cases}\dot{x}=-\sigma x+\sigma y,&x(0)=x_{0}\\ \dot{y}=-\sigma x-y-xZ,&y(0)=y_{0}\\ \dot{Z}=-bZ+XY-b(r+\sigma),&Z(0)=Z_{0}.\end{cases}$ (4)
In the above systems, Equations˜2, 3 and 4, correspond to the regimes when the observed values of $X$, $Y$, or $Z$, respectively, are observed and inserted directly into the equations for the remaining, unobserved components. That is, e.g. in Equation˜2 we observe the true value of $X$ continuously in time and insert that into the evolution equations of $y$ and $z$. In these equations we note that numerically we should not be simulating the evolution of $X$, $Y$, or $Z$ as they appear in Equations˜2, 3 and 4, respectively, but rather these values are observed (given) from the true solution in Equation˜1 and are inserted directly into the equations for the remaining variables. Analysis on Equations˜2, 3 and 4 was conducted in [47] where it was shown that the continuous coupling on $X$ or $Y$ alone was sufficient to obtain synchronization between the assimilated equations and the reference solution, however the coupling seen in Equation˜4 was not sufficient to obtain convergence in either the $x$ or $y$ components to the corresponding components $X$ or $Y$ of the true solution. We do note that in [47] the equations for Equation˜1 were written in a slightly different form, however this can be easily reconciled by using the transformation $\tilde{Z}=Z+(r+\sigma)$.
For the sake of completeness, we will discuss the convergence results for the above systems, Equations˜2, 3 and 4 proven in [47] and discuss how the results break down in the absence of dissipation. Let
$\delta=(X-x,Y-y,Z-z),$ (5)
where $(X,Y,Z)$ is the given solution of the original Lorenz system Equation˜1 and $(x,y,z)$ is the solution to one of the assimilated equation, Equations˜2, 3 and 4. Here we define $\delta$ with the understanding that the variable in $(x,y,z)$ missing from the assimilated equation is taken to be equal to the corresponding variable in Equation˜1, e.g. when considering Equation˜2 we define $x(t)\mathrel{\mathop{\ordinarycolon}}=X(t)$ for all time, $t$ and have $\delta=(0,Y-y,Z-z)$. The convergence results for the synchronization algorithms, Equations˜2, 3 and 4 are given as follows:
Theorem 2.1 (Hayden (2007) [47]).
Let $(X,Y,Z)$, be the solution to Equation˜1, then the following are true when $\sigma>0,b>0,$ and $r>0$.
If $(x,y,z)$ is the solution to Equation˜2, then
$\mathinner{\!\left\lVert\delta(t)\right\rVert}^{2}_{\ell^{2}}\leq\mathinner{\!\left\lVert\delta_{0}\right\rVert}_{\ell^{2}}e^{-2t}.$ (6)
If $(x,y,z)$ is the solution to Equation˜3, then
$\begin{split}\mathinner{\!\left\lVert\delta(t)\right\rVert}^{2}_{\ell^{2}}&\leq|X(0)-x(0)|e^{-2\sigma}+|Z(0)-z(0)|e^{-bt}\\ &\quad+\frac{J|X(0)-x(0)|^{2}}{b(2\sigma-b)}\left(e^{-bt}-e^{-2\sigma t}\right),\end{split}$ (7)
where
$J=\begin{cases}r^{2},&b\leq 2\\ \frac{b^{2}r^{2}}{4(b-1)},&b\geq 2\end{cases}$ (8)
The proofs of Theorem˜2.1 are derived by taking the difference of the assimilated equation (Equation˜2 or Equation˜3) with Equation˜1, summing the squares of the derivatives, and applying Grönwall’s inequality to the resulting quantity. The $J$ estimate in Equation˜7 arises from the use of the bound
$\limsup_{t\to\infty}\left(Y(t)\right)^{2}\leq J,$ (9)
which was derived in [27] and was used to bound the contribution from the nonlinear term in the $z$ equation. It is worth noting that the bound Equation˜9 is valid uniformly and globally in time for solutions in the global attractor that we intend to recover using data assimilation algorithms.
We note that the analysis done in the proof of Theorem˜2.1 is dependent on the particular parameters and requires $\sigma>0$, $b>0$ and $r>0$. To illustrate how these sorts of estimates break down in the absence of dissipation, we consider the case when both $X$ and $Y$ are observed, i.e., are given from the true solution of Equation˜1. This leads to the following system of equations
$\begin{cases}\dot{X}=-\sigma X+\sigma Y,&X(0)=X_{0}\\ \dot{Y}=-\sigma X-Y-XZ,&Y(0)=Y_{0}\\ \dot{z}=-bz+XY-b(r+\sigma),&z(0)=z_{0}.\end{cases}$ (10)
Now, if we take w=zZw=z-Z, where $Z$ solves Equation˜1, we see from Equations˜1 and 10 that $w$ satisfies the following equation
$\dot{w}=-bw,\quad w(0)=z_{0}-Z_{0}.$ (11)
$|w(t)|^{2}\leq|w(0)|^{2}e^{-bt},$ (12)
which yields convergence of $z$ to $Z$ exponentially fast in time. If we instead consider $b=0$, removing the dissipation on the $Z$ component of Equation˜1, then we now have that $\dot{w}=0$, and so $w(t)=z_{0}-Z_{0}$ holds for all time. Therefore, we have that the absence of dissipation on the $Z$ component prevents synchronization between the assimilated solution and the reference solution, even when both the $X$ and $Y$ components are observed continuously in time.
We note that the above discussion has been on the synchronization data assimilation scheme, given by the direct insertion, as it was first introduced in [67], of the observed quantities into the model equations. We now consider the nudged Lorenz system for the observed $X$ and $Y$ components of Equation˜1, given as follows:
$\begin{cases}\dot{x}=-\sigma x+\sigma y+\mu(X-x)\\ \dot{y}=-\sigma x-y-xz+\mu(Y-y)\\ \dot{z}=-bz+xy-b(r+\sigma)\end{cases}$ (13)
At any fixed time, $t$, the solutions, $(x(t;\mu),y(t;\mu),z(t;\mu))$, to the nudged system, Equation˜13, are expected to converge to the solution $(X(t),Y(t),z(t))$ of the synchronization algorithm, given by Equation˜10 as $\mu\to\infty$. This can be justified rigorously by applying the argument present in [15] to the Lorenz 1963 system, provided both methods are initialized with the same initial data. An analogous statement was made in [71] for the Lorenz system in particular (with $b>0$), where it was shown that the solution, $x(t;\mu)$, of Equation˜13, converges to $X(t)$ at any fixed time, $t$, as we send $\mu\to\infty$ when the nudging term is present only on the equation for $x$ in Equation˜13. In the case where $b=0$, if we assume a priori that the solutions of Equation˜13 converge to those of Equation˜10 as $\mu\to\infty$, one can deduce using the reverse triangle inequality, that $\mathinner{\!\left\lVert\delta(t)\right\rVert}^{2}_{\ell^{2}}\geq|z_{0}-Z_{0}|^{2}$, where $\delta$ is given as defined previously in Equation˜5. Therefore, we expect the nudged Lorenz system, Equation˜13, to fail to recover the true solution in the same way that Equation˜10 fails when dissipation is removed from the equations for $z$.
We now demonstrate numerically that this dissipation mechanism (when $\sigma,b>0$) of the Lorenz system is necessary for data assimilation methods to work. In particular, we will be considering a partially dissipative variant of the Lorenz system by taking $b=0$, i.e., eliminating the dissipation mechanism only on the $Z$ component. The analytical results regarding the nudging algorithm applied to the Lorenz equations have been done when $b>0$ and are qualitatively similar to those we reference above in Theorem˜2.1 from [47], see e.g. [11, 71, 17]. That is, the solutions of Equation˜13 synchronize with those of Equation˜1 exponentially fast in time for large enough nudging parameter values and with the exponential rate depending on the particular parameters of Equation˜1. While nudging is only required on either the $X$ or $Y$ components in order to drive convergence, we will utilize nudging on both the $X$ and $Y$ components to emphasize that, even with additional nudging, the method is not effective when the dissipation on the $z$-component is eliminated (i.e. when $b=0$).
First, we verify the validity of nudging on the Lorenz system with full dissipation. Using the initial data and parameter choices discussed previously we apply nudging to the Lorenz system with $b=8/3$. The results can be seen in Figure˜1, where we have applied nudging to only the first two components of Equation˜13. Here we can see that all variables converge to an error level of approximately 1e121e-12 with spurious dips below this error level in each component. Note that here we see exponential convergence to the true solution, as we expect in this case.
Next, we examine what happens when we apply nudging for the partially dissipative Lorenz system, i.e., with $b=0$. The results can be seen in Figure˜2. Note that here we still see exponential decay to the true solution in both $x$ and $y$. While nudging appears to affect $z$ initially, this effect quickly fades after a short interval of time and the error appears to stabilize to be approximately constant $1e+1$. The error level that is attained appears to be directly linked to the difference in the initial conditions for the $z$ component. To illustrate this, we plotted the $\ell^{2}$ (Euclidean norm) error over time when both the reference and nudged equations are initialized with the same initial data, except for the $z$ component. Here $Z(t_{0})=50$ and $z(t_{0})=50+\delta$, where $\delta$ ranges in values from 1e11e-1 to 1e121e-12. The results of these tests can be seen in Figure˜3 and they illustrate that the $\ell^{2}$ error is highly sensitive to small perturbations in the initial condition for $z$. We note that the behavior seen in Figure˜3 appears to align with what one should expect from our earlier discussion of Equation˜11 for the synchronization algorithm. We note that we obtain an initial exponential decay in the error in the $z$ component which is driven by the convergence in both $x$ and $y$. This initial convergence levels off and becomes approximately constant over time.
3. 1D PDEs - Korteweg-de Vries
3.1. Preliminaries and Analytical Results
@@
Generally speaking, the nudging algorithm is described as follows. Consider the following evolution system:
$u_{t}=F(u),$ (14)
@@
where here $F$ is a given differential operator that is potentially nonlocal and nonlinear. We will specifically be considering solutions of Equation˜14 that lie in, e.g., the phase space $L^{2}$ subject to periodic boundary conditions.
@@
In the remainder of this section we will utilize $H^{m}_{\text{per}}$, with $m$ being a non-negative integer, to denote the closure of the set of $L$-periodic trigonometric polynomials on $\mathbb{R}$, with respect to the standard Hilbert norm:
$\mathinner{\!\left\lVert f\right\rVert}_{H^{m}_{\text{per}}}\mathrel{\mathop{\ordinarycolon}}=\left(\sum_{k=0}^{m}\int_{0}^{L}\mathinner{\!\left\lvert\frac{\partial^{k}f}{\partial x^{k}}\right\rvert}^{2}dx\right)^{\frac{1}{2}}.$ (15)
@@
Definition 3.1 (Determining modes, as given in [37, 54, 23, 36]).
We say that Equation˜14 has finitely many determining modes if there exists a positive integer $M>0$ such that for any two given solutions $u_{1}$ and $u_{2}$ of Equation˜14 the following holds:
@@
$\lim_{t\to\infty}\mathinner{\!\left\lVert P_{M}(u_{1}(t)-u_{2}(t))\right\rVert}_{L^{2}}=0,$ (16)
@@
$\lim_{t\to\infty}\mathinner{\!\left\lVert u_{1}(t)-u_{2}(t)\right\rVert}_{L^{2}}=0.$ (17)
@@
We assume that Equation˜14 is globally well-posed and for such systems we define the related nudged system, introduced in [4, 5], as follows:
$v_{t}=F(v)+\mu I_{h}(u-v).$ (18)
Here $I_{h}$ is a linear spatial interpolant with associated length scale $h$ and $\mu$ is a constant feedback-control nudging parameter. For our purposes we will consider only $I_{h}\mathrel{\mathop{\ordinarycolon}}=P_{M}$, where $P_{M}$ is given to be a projection operator onto the lowest $M$ Fourier modes, i.e. $h\sim\frac{L}{M}$, where $L$ is the period for the underlying evolutionary equation Equation˜14. As we require $P_{M}u$ to be well defined we will only consider solutions to Equation˜14 that are in $L^{2}_{\text{per}}$ for all positive time.
@@
$u_{t}+uu_{x}+\delta^{2}u_{xxx}=0.$ (19)
Here $u\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\times[0,\infty)\to\mathbb{R}$ represents the amplitude of a wave and $\delta$ is the dispersive coefficient. We consider the KdV equation equipped with the periodic boundary condition $u(x,t)=u(x+2,t)$, for all $x\in\mathbb{R}$ and all time $t\geq 0$ with spatial mean zero. Hence, the period in this case is $L=2$. We note that choosing the length of the spatial domain to be 2 is arbitrary and was chosen as this choice of domain appears in the literature, see, e.g., [80].
@@
$v_{t}+vv_{x}+\delta^{2}v_{xxx}=\mu P_{M}(u-v).$ (20)
Here $\mu>0$ is a feedback-control nudging parameter and $P_{M}$ is given to be the projection operator onto the lowest $M$ Fourier modes. Here $u$ is the unknown reference solution of Equation˜19, for which the first $M$ Fourier modes are observed (given) and we aim to reproduce it.
@@
$\hat{f}(k,t)\mathrel{\mathop{\ordinarycolon}}=\int_{0}^{2}f(x,t)e^{-\pi ik\cdot x}dx.$ (21)
For simplicity of presentation, we omit the time index and, unless otherwise stated, we write $\hat{f}_{k}\mathrel{\mathop{\ordinarycolon}}=\hat{f}(k,t)$.
@@
Proposition 3.2.
Let $k\in\mathbb{Z}\setminus\{0\}$ and $\phi\in L^{2}_{\text{per}}$ with $\phi$ having spatial mean zero. The following hold true:
@@
    $\phi(x+\frac{L}{k})=\phi(x)$ for a.e. $x\in\mathbb{R}$ if and only if
    $\phi(x)=\sum_{n\in k\mathbb{Z}}\hat{\phi}_{n}e^{i\frac{2\pi}{L}nx}.$ (22)
@@
    Let $M$ be a positive integer and let $k$ be an integer satisfying $\mathinner{\!\left\lvert k\right\rvert}\geq M$. Suppose $\phi(x+\frac{L}{k})=\phi(x)$ for a.e. $x\in\mathbb{R}$, then
    $P_{M}\phi=0$ (23)
@@
Proposition 3.3.
Suppose $k\in\mathbb{Z}\setminus\{0\}$ and $u^{in},f\in H^{2}_{\text{per}}$ such that $f(x+\frac{L}{k})=f(x)$ and $u^{in}(x+\frac{L}{k})=u^{in}(x)$ for all $x\in\mathbb{R}$ and $u^{in}$ and $f$ both have spatial mean zero. Let $u(x,t)$ be the corresponding solution of the damped and driven KdV equation:
$u_{t}+uu_{x}+\delta^{2}u_{xxx}+\gamma u=\beta f,\\ u(x,0)=u^{in}(x),$ (24)
with $\gamma\geq 0$ and $\mathinner{\!\left\lvert\beta\right\rvert}\geq 0$, subject to periodic boundary conditions with period $L$.
Then $u(x+\frac{L}{k},t)=u(x,t)$ for a.e. $x\in\mathbb{R}$ and all $t\in\mathbb{R}$. Moreover, for every $0<T<\infty$, the solution $u$, restricted to the time interval [T,T][-T,T], is continuous and bounded, as in [12].